# Phase behavior of Cacio and Pepe sauce

### v2

@@ IV A minimal model recapitulates the mixture phase behavior
Here, we introduce a minimal model that qualitatively matches the behavior of the mixture when the starch concentration is kept constant. Choosing which components to explicitly include in the theoretical description is a delicate step. This is because our system contains starch, salt, and lipids, together with two different kinds of proteins in the cheese, namely casein and whey (see also Fig. 4 a). Furthermore, in the range of temperatures explored, a fraction of whey proteins undergoes denaturation [29, 30]. In the following, we assume that the shape of the binodal line in the phase diagram Fig. 3 c) can be described by the phase separation of a binary mixture. We chose the binary mixture framework for its simplicity, but the applicability of such an approach has surely many limitations. In the final discussion, we outline more realistic modeling approaches and their drawbacks.
As components of our model, we chose the cheese proteins (whey and casein) and an effective solvent encompassing water, starch, salt, and lipids constituting the cheese. We indicate the mass fractions of proteins and solvent with $\phi$, and $\phi_{\text{s}}$, respectively. Mass conservation implies $\phi+\phi_{\text{s}}=1$. If both the components have almost equal densities, the phase diagram of a mixture in terms of the mass fraction of their components can be derived from the following free energy density
@@
These equations have to be simultaneously solved at each temperature to determine the protein mass fraction in the coexisting phases, $\phi^{\text{I}}$ and $\phi^{\text{II}}$. As a function of temperature, $\phi^{\text{I}}$ and $\phi^{\text{II}}$ span the binodal line observed in Fig. 3 c). Within the framework of binary mixtures, the minimum of the paraboloid-shaped binodal corresponds to a lower critical solution temperature. At this critical temperature, slightly above $60\degree\text{C}$, the two solutions collapse, i.e. $\phi^{\text{I}}=\phi^{\text{II}}$. In principle, one has to find the right parameters $n$ and $\chi$ so that the binodal curve we observe is duly reconstructed by this minimal model. Notice that inferring the dependence of these parameters on temperature can be challenging. For this reason, we employ a reverse-engineering procedure, in analogy with [32]. We analytically solve Eq. (1) for $\chi$ and $n$ for each value of the temperature $T$, obtaining:
$\chi(T)=\mathcal{F}(\phi^{\text{I}},\phi^{\text{II}})\qquad n(T)=\mathcal{G}(\phi^{\text{I}},\phi^{\text{II}})$ (2)
@@
where $\mathcal{F}$ and $\mathcal{G}$ are specified in Eq. S1. Here, $\phi^{\text{I}}$ and $\phi^{\text{II}}$ can be directly extracted, for each value of $T$, from the experimental binodal in Fig. 3 c). The resulting dependence on temperature of $\chi(T)$ (in red) and $n(T)$ (in blue) is reported in Fig. 4 b) and constitutes the only solution for these two parameters compatible with experiments. From a microscopic point of view, an increase in the interaction parameter $\chi$ can be due to the fact that heat induces denaturation of whey proteins, with consequential aggregation, while simultaneously favoring whey-micelle and micelle-micelle interactions [27, 28, 33] (see Fig. 4 a)). On the other hand, casein micelles are relatively heat stable, undergoing negligible dissociation on heating [34]. On the same line, the increase in the relative size might be associated with whey denaturation, or it might be due to the fact that micelle and whey proteins can form starch-mediated complexes that have a larger size [35]. Furthermore, in Fig. 4 c) we reconstruct the binodal line employing a direct approach to Eq. (1) using the linear fits of $\chi(T)$ and $n(T)$ as parameters. Finally, notice that starch polymers are roughly $100$ times larger than proteins while being $100$ times less abundant in solution. As a consequence, we expected a relative size of the order of (or slightly greater than) the unity. This expectation is confirmed by our analysis, hinting at the fact that our model, despite its simplicity, is able to capture the main features of the observed phase behaviour.
Figure 4: A minimal model for the phase behavior of Cacio and pepe sauce. a) Cheese is composed of casein organized in micelles together with calcium phosphate and a small percentage of whey proteins. Upon heating, whey proteins denaturate, reaching a state that favors whey-whey and whey-casein aggregation. Furthermore, casein micelles aggregate on heating. b) Interaction and relative size (respectively $\chi$ and $n$ in the effective free energy) obtained from experiments by solving Eq. (1) for different temperatures. c) Comparison between theoretical and experimental binodal curve.
V Scientific recipe
@@
A true Italian grandmother or a skilled home chef from Rome would never need a scientific recipe for Cacio and pepe, relying instead on instinct and years of experience. For everyone else, this guide offers a practical way to master the dish. Preparing Cacio and pepe successfully depends on getting the balance just right, particularly the ratio of starch to cheese. The concentration of starch plays a crucial role in keeping the sauce creamy and smooth, without clumps or separation. If the starch content is less than 1% of the cheese weight, the sauce is prone to separating into unpleasant system-sized clumps corresponding to the “mozzarella phase” in Figures 1 and 2. On the other hand, exceeding 4% starch results in a sauce that becomes stiff and unappetizing as it cools. The ideal range, as confirmed by both taste and texture tests, lies between 2% and 3%, ensuring stability and a pleasant consistency.
@@
For a practical example, consider preparing Cacio and pepe for two hungry people. This typically requires 240 grams of pasta (tonnarelli is preferred, though spaghetti or rigatoni also works well) and 160 grams of cheese. Traditionalists would insist on using only pecorino, but some argue that up to 30% Parmesan is acceptable, though this remains a point of debate. To achieve the correct starch ratio, 4 grams of starch is optimal for 160 grams of cheese.
@@
The pasta water alone does not contain enough starch to stabilize the sauce effectively. As we already discussed, one could use pasta water “risottata”, i.e. boiled down to concentrate the starch, but the process offers little control over the final starch amount. A more precise and reliable method is to dissolve 4 grams of powdered starch (such as potato or corn starch) in 40 grams of water. Heat this mixture gently until it thickens and turns from cloudy to nearly clear. This process, known as starch gelation, increases the viscosity of the mixture. Once the starch solution is ready, allow it to cool slightly by either waiting or mixing it with a small amount of cold water. The next step is to combine the starch solution with the cheese. Manually grating the cheese is not ideal, since it leads to chunks of different sizes. We recommend blending it with the starch solution for a smooth, homogeneous sauce. If blending is difficult, add a splash of water to ease the process, and season with freshly ground black pepper to taste.
@@
Meanwhile, cook the pasta in slightly salted water until it is al dente. Save some of the pasta cooking water before draining. Once the pasta has been drained, let it cool down for up to a minute (even a little bit longer for an amount of pasta >1kg) to prevent the heat from destabilizing the sauce. Mix the pasta with the sauce, ensuring even coating, and adjust the consistency by gradually adding reserved pasta water as needed.
@@
One of the benefits of this stabilized sauce is its ability to withstand reheating. Unlike traditional methods that risk clumping or separation, this sauce maintains its texture and stability even when brought to temperatures in the order of $80-90\degree\text{C}$. This ensures the dish can be served hot, allowing diners to enjoy it at its best.
@@
This method offers a simple yet precise way to consistently achieve a perfect Cacio and pepe. This recipe is inspired by Luciano Monosilio’s YouTube video [36], though it does not include olive oil as suggested in his version. Despite this difference, both recipes share a focus on respecting the tradition while ensuring a reliable and enjoyable result.
VI Discussion
@@
In this work, we investigated the phase behavior of Cacio and pepe, one of most famous and complicated pasta recipes of Italian culinary tradition. We quantified the stabilizing role of the starch when mixed with cheese and water, which favors the homogeneity of the mixture up to temperatures in the order of water boiling point. Also, by inspecting the role of cheese protein concentration, we unveil an unforeseen binodal curve resembling the one of a phase separating system with a lower critical solution temperature. By employing a minimal model to capture this phenomenology, we were able to extrapolate how protein-solution interaction and relative size behave as a function of temperature. Although it is difficult to map exactly our observations to the microscopics of this complex system, we rationalized these effective phenomena, shedding light on the interplay between cheese protein and starch under heating conditions, a widespread scenario in culinary experiences. Ultimately, our approach leads to the formulation of a scientific recipe for Cacio and pepe that capitalizes on our findings and highlights their applicative perspective.
A potential future direction could be to better explain the starch-dependent morphology of the dense phase. Indeed, as noticed in Figs. 1 and 2, the cheese-rich phase abruptly switches from a big clump (that we named “Mozzarella phase” ) to many small clusters, as starch concentration increases. This phenomenon could be explained by using a model that explicitly considers at least three different species.
@@
The simplest description would account for proteins in two states A and B with stronger and weaker interaction propensity, respectively, and an effective solvent encompassing water and all remaining molecules. Then one would need to account for transitions between the two states of the proteins, controlled by temperature and starch amount. In the spirit of Ref. [37], the mozzarella-like clump could represent a phase where A proteins are dominant, while small clusters could arise whenever B proteins drive phase separation. From a microscopic point of view, A proteins could correspond to denatured whey proteins, while B proteins could encompass folded whey proteins and unfolded ones sequestered by the starch. One could even include casein micelles explicitly as a fourth component, instead of incorporating them in the effective solvent. An even more detailed model could consider protein aggregates of different sizes explicitly. One then could couple phase separation and gelation, describing the switch to a mozzarella phase as the onset of a gel phase [38, 39, 40]. Such generalizations of our model could shed more light on the molecular features of this system, eventually obtaining novel insights not only into perfecting the recipe but also in the broad field of food science. However, the main issue with these multi-component approaches is that they require more experiments to derive the dependencies of all interaction and size parameters on temperature. For example, it would be necessary to quantify the fraction of denatured whey protein for each temperature in each phase, which is a hard task.
@@
Other interesting future directions might involve a more in-depth analysis of how starch affects effective parameters and influences the viscosity of Cacio and pepe sauce, and the potential role of pepper grains, another important part of the original recipe, as aggregation nuclei. We hope this paper ignited the idea that the genuine passion for fine cuisine can be translated into insightful results to refine complex preparations and make them achievable with accessible kitchen tools.
@@ Acknowledgments
The authors thank Frank Jülicher, for supporting the initiative and stimulating discussions. We thank Tetsuya Spippayashi for enlightening clarifications on the historical origins of Cacio and pepe and Giuseppe Ricchitelli for helping with the construction of the experimental apparatus. We further thank Martina Gaiba, Alessandro Gaiba, John D. Treado, Virginia Lepore, Eleonora Nanu, Julia Kirsch, Lara Koehler, Burak Budanur, Elizabeth Brückner, and M.J. Franco Oñate for their support and for eating up the sample leftovers. Credit for Figure 1a) is attributed to unastellaincucina.
@@
Supplemental Material for
Phase separation in cacio e pepe
S1 Equipment and experimental protocol
@@
    Sauce preparation: the ingredients for a small batch of sauce ($\approx 110$ g, with the exact weight depending on the specific formulation) are mixed in a clean mug to achieve the desired composition of cacio cheese, starch, and water. The mixture is homogenized using an immersion blender, then transferred into a saucepan and weighed. The saucepan is then placed into the wooden platform of the heating apparatus (Figure S1a).
@@
    Temperature ramp: the sauce is gradually heated up while its actual temperature of is continuously monitored using a digital cooking thermometer with a probe immersed in the sauce itself. During the heating process, the mixture is constantly stirred with a spoon to ensure uniform heating and to prevent the formation of large cheese aggregates on the saucepan walls.
@@
    Image acquisition: the sauce sample withdrawn from the saucepan is quickly transferred into a clean Petri dish. The mixture is evenly spread across the dish by gently shaking it, and then the dish is positioned on the transparent support (Figure S1b). A photograph of the sample is captured using the cell phone. After imaging, the sauce sample is collected in a separate container for later consumption. This entire operation is optimized to be completed within approximately 20 seconds, ensuring that no significant precipitation of cheese clumps occurs during the sample’s transfer and photography.
@@
Figure S3: Three examples of segmented images illustrating varying degrees of phase separation: a, b) low, c, d) medium, and e, f) high. Panels a), c), and e) show cropped grayscale experimental images (not to scale), where darker regions represent aggregates. Corresponding segmented aggregates are displayed in panels b), d), and f), obtained using the histogram-based segmentation method described below.
The data analysis involved segmenting aggregates in petri dish images and computing their elongation. Initially, the images were converted to greyscale and normalized to an intensity range of $[0,1]$. To enhance visibility, we adjusted exposure using the equalize_adapthist function from the exposure module in scikit-image, with a clip limit of $0.01$. Images were then smoothed with a Gaussian filter using a sigma of $1$ pixel. These parameters were determined based on iterative trials.
@@
The segmented aggregates were then analyzed using the regionprops routine from scikit-image. In particular, we reported the mean aggregate elongation in the phase diagrams, calculated using the major axis length property from regionprops.
To give an idea of how elongation is computed, one first calculates the covariance matrix $C_{ab}$ for the pixel coordinates $\{x_{i},y_{i}\}$ of each aggregate, using the following flat measure:
@@
with analogous formulas for $C_{xx}$ and $C_{yy}$. A covariance matrix carries the information of the aggregate area as $\sqrt{\det C}$. However, this area differs from the actual aggregate area, and so the matrix $C_{ab}$ was scaled such that $\det C=A^{2}$, where $A$ is the area of the segmented region. The largest eigenvalue of the rescaled $C_{ab}$ corresponds to the square of the aggregate’s elongation.

### v3

@@ IV A minimal model recapitulates the mixture phase behavior
Here, we introduce a minimal model that qualitatively matches the behavior of the mixture when the starch concentration is kept constant. Choosing which components to explicitly include in the theoretical description is a delicate step. This is because our system contains starch, salt, and lipids, together with two different kinds of proteins in the cheese, namely casein and whey (see also Fig. 4(a)). Furthermore, in the range of temperatures explored, a fraction of whey proteins undergoes denaturation [38, 39]. In the following, we assume that the shape of the binodal line in the phase diagram in Fig. 3(c) can be described by the phase separation of a binary mixture. We chose the binary mixture framework for its simplicity, but the applicability of such an approach has surely many limitations. In the final discussion, we outline more realistic modeling approaches and their drawbacks.
As components of our model, as mentioned in the previous section, we chose the cheese proteins (whey and casein) and an effective solvent encompassing water, starch, salt, and lipids constituting the cheese. We indicate the mass fractions of proteins and solvent with $\phi$, and $\phi_{\text{s}}$, respectively. Mass conservation implies $\phi+\phi_{\text{s}}=1$. If both the components have almost equal densities, the phase diagram of a mixture in terms of the mass fraction of their components can be derived from the following free energy density [40, 41]
@@
These equations have to be simultaneously solved at each temperature to determine the protein mass fraction in the coexisting phases, $\phi^{\text{I}}$ and $\phi^{\text{II}}$. As a function of temperature, $\phi^{\text{I}}$ and $\phi^{\text{II}}$ span the binodal line observed in Fig. 3(c). Within the framework of binary mixtures, the minimum of the paraboloid-shaped binodal corresponds to a lower critical solution temperature. At this critical temperature, slightly above $60\degree\text{C}$, the two solutions collapse, i.e. $\phi^{\text{I}}=\phi^{\text{II}}$. In principle, one has to find the right parameters $n$ and $\chi$ so that the binodal curve we observe is duly reconstructed by this minimal model. Notice that inferring the dependence of these parameters on temperature can be challenging. For this reason, we employ a reverse-engineering procedure, in analogy with [43]. We analytically solve Eq. (1) for $\chi$ and $n$ for each value of the temperature $T$, obtaining:
$\chi(T)=\mathcal{F}(\phi^{\text{I}},\phi^{\text{II}})\qquad n(T)=\mathcal{G}(\phi^{\text{I}},\phi^{\text{II}})\,,$ (2)
@@
where $\mathcal{F}$ and $\mathcal{G}$ are specified in Eq. (S1) of the supplementary material. Here, $\phi^{\text{I}}$ and $\phi^{\text{II}}$ can be directly extracted, for each value of $T$, from the experimental binodal in Fig. 3(c). The resulting dependence on temperature of $\chi(T)$ (in red) and $n(T)$ (in blue) is reported in Fig. 4(b) and constitutes the only solution for these two parameters compatible with experiments. From a microscopic point of view, an increase in the interaction parameter $\chi$ can be due to the fact that heat induces denaturation of whey proteins, with consequential aggregation, while simultaneously favoring whey-micelle and micelle-micelle interactions [34, 35, 44] (see Fig. 4(a)). On the other hand, casein micelles are relatively heat stable, undergoing negligible dissociation on heating [45]. On the same line, the increase in the relative size might be associated with whey denaturation, or it might be due to the fact that micelle and whey proteins can form starch-mediated complexes that have a larger size [46]. Furthermore, in Fig. 4(c) we reconstruct the binodal line employing a direct approach to Eq. (1) using the linear fits of $\chi(T)$ and $n(T)$ as parameters. Finally, notice that starch polymers are roughly $100$ times larger than proteins while being $100$ times less abundant in solution. As a consequence, we expected a relative size of the order of (or slightly greater than) the unity. This expectation is confirmed by our analysis, hinting at the fact that our model, despite its simplicity, is able to capture the main features of the observed phase behaviour.
Figure 5: Effect of trisodium citrate on the stability of Cacio e pepe sauce. The diagram shows the behavior of the sauce as a function of temperature for different concentrations of sodium citrate (1%, 2%, and 3% relative to cheese mass) at a fixed 1:1 cheese-to-water ratio. Below 2% citrate, the sauce remains unstable, leading to system-wide aggregation characteristic of the Mozzarella Phase. For concentrations of 2% and above, the sauce is fully stabilized, preventing any visible aggregation.
V Trisodium Citrate as an Alternative Stabilizer
In this section, we investigate the effects of trisodium citrate (Na3C6H5O7) on the stability of the sauce. Trisodium citrate is widely used in the food industry as an emulsifier and pH buffer. Its ability to chelate calcium ions and alter protein interactions makes it particularly effective in stabilizing emulsions and improving the texture of dairy-based sauces. In particular, its chelating action primarily affects casein proteins, as calcium plays a key role in their aggregation. It is commonly used in processed cheese production to enhance smoothness and prevent undesired aggregation.
To quantify its effect in the context of Cacio e pepe, we applied our experimental protocol to a mixture of cheese and water at a 1:1 ratio, adding 1%, 2%, and 3% of trisodium citrate relative to cheese mass. The results of our experiments are shown in Fig. 5. We observe that at concentrations below 2%, the citrate is insufficient to stabilize the sauce, leading to the formation of system-wide clumps characteristic of the Mozzarella Phase, albeit at higher temperatures compared to the sauce without any stabilizer (see Fig. 5, second vs. first column). While these clumps qualitatively differ from those in the absence of citrate in terms of consistency, they still span the entire system.
For concentrations of 2% and above, however, the sauce remains highly stable, with no visible aggregates of any size forming. This highlights that the transition to a fully stable sauce is sharper with trisodium citrate compared to starch, where small aggregates persist even in the regime where the Mozzarella Phase is absent. This difference is expected, as the macroscopic behavior directly reflects the distinct molecular mechanisms at play when citrate is present. These observations suggest that the latter can serve as an effective alternative stabilizer for Cacio e pepe, offering a reliable way to control sauce consistency, though at the cost of deviating from strict culinary tradition. However, while the sauce stabilization is more efficient, we found the taste of the cheese to be slightly blunted, likely due to the basic properties of the salt.
VI Scientific recipe
@@
A true Italian grandmother or a skilled home chef from Rome would never need a scientific recipe for Cacio e pepe, relying instead on instinct and years of experience. For everyone else, this guide offers a practical way to master the dish. Preparing Cacio e pepe successfully depends on getting the balance just right, particularly the ratio of starch to cheese. The concentration of starch plays a crucial role in keeping the sauce creamy and smooth, without clumps or separation. If the starch content is less than 1% of the cheese weight, the sauce is prone to separating into unpleasant system-sized clumps corresponding to the Mozzarella Phase in Figs. 1 and 2. On the other hand, exceeding 4% of starch results in a sauce that becomes stiff and unappetizing as it cools. The ideal range, as confirmed by both taste and texture tests, lies between 2% and 3%, ensuring stability and a pleasant consistency.
@@
For a practical example, consider preparing Cacio e pepe for two hungry people. This typically requires 300g of pasta (tonnarelli is preferred, though spaghetti or rigatoni also works well) and 200g of cheese. The amount of cheese can, of course, vary depending on personal taste. Traditionalists would insist on using only Pecorino Romano DOP, but some argue that up to 30% Parmigiano Reggiano DOP is acceptable, though this remains a point of debate. To achieve the correct starch ratio, 5g of starch is optimal for 200g of cheese.
@@
The pasta water alone does not contain enough starch to stabilize the sauce effectively. As we already discussed, one could use pasta water “risottata”, i.e., boiled down to concentrate the starch, but the process offers little control over the final starch amount. A more precise and reliable method is to dissolve 5g of powdered starch (such as potato or corn starch) in 50g of water. Heat this mixture gently until it thickens and turns from cloudy to nearly clear. This transition, known as starch gelatinization, is characterized by a sudden increase in viscosity. The next step is to combine the starch gel with the cheese. Manually grating the cheese is not ideal, since it may lead to chunks of different sizes. We recommend blending it with the starch solution for a smooth, homogeneous sauce. To ease blending and also cool the starch gel, add 100g of water to it before adding the cheese. This increases the water-to-cheese ratio to 75%, where, as suggested by Fig. 3, the sauce is even more stable than the sauce with the same amount of starch but at 100% water content shown in Fig. 2. Finish the sauce by adding the black pepper to the mixture. To enhance its flavors and aromas, we recommend toasting it in a pan beforehand.
Alternatively, one could replace starch with 5g of trisodium citrate (2.5% of cheese mass, dissolved in 150g of water), which provides excellent stabilization properties. However, this substitution comes at the cost of tradition, as gelatinized starch is technically an ingredient of the original recipe, while citrate is not. Also, as discussed in the previous section, while citrate stabilizes the sauce more effectively, it slightly blunts the cheese’s flavor, likely because of its basic properties, and changes the mouthfeel of the sauce, though not necessarily in a negative way.
@@
Meanwhile, cook the pasta in slightly salted water until it is al dente. Save some of the pasta cooking water before draining. Once the pasta has been drained, let it cool down for up to a minute (even a little bit longer for an amount of pasta $\geq$1kg) to prevent the excessive heat from destabilizing the sauce. Indeed, in our experiments, we gradually increased the temperature of the sample over approximately 40 minutes (from 50C to 95C) to minimize kinetic effects, leading to the results shown in Fig. 2 and Fig. 3. While this level of control is impractical in a kitchen setting, it is important to limit thermal shocks as much as possible to achieve results closer to those observed in our experiments. Finally, mix the pasta with the sauce, ensuring even coating, and adjust the consistency by gradually adding reserved pasta water as needed.
@@
One of the benefits of this stabilized sauce is its ability to withstand reheating, most effectively in a pan. Unlike traditional methods that risk clumping or separation, this sauce maintains its texture and stability even when brought to temperatures in the order of $80-90\degree\text{C}$. This ensures the dish can be served hot, allowing diners to enjoy it at its best. Garnish with grated cheese and pepper before serving.
@@
This method offers a simple yet precise way to consistently achieve a perfect Cacio e pepe. It is particularly useful for cooking large batches of pasta, where heat control can be challenging and requires extra care. This recipe is inspired by Luciano Monosilio’s YouTube video [47], though it does not include olive oil as suggested in his version. Despite this difference, both recipes share a focus on respecting tradition while ensuring a reliable and enjoyable result.
VII Discussion
@@
In this work, we investigated the phase behavior of Cacio e pepe, one of the most famous and complicated pasta recipes of Italian culinary tradition. We quantified the stabilizing role of starch when mixed with cheese and water, which favors the homogeneity of the mixture up to temperatures in the order of the water boiling point. Also, by inspecting the role of cheese protein concentration, we unveil an unforeseen binodal curve resembling that of a phase-separating system with a lower critical solution temperature. By employing a minimal model to capture this phenomenology, we were able to extrapolate how protein-solution interaction and relative size behave as a function of temperature. Although it is difficult to map exactly our observations to the microscopics of this complex system, we rationalized these effective phenomena, shedding light on the interplay between cheese protein and starch under heating conditions, a widespread scenario in culinary experiences. Additionally, we explored the stabilizing effect of trisodium citrate, which, unlike starch, induces a sharp transition from a phase-separated sauce to a homogeneous emulsion by chelating calcium ions and preventing protein aggregation. Ultimately, our approach leads to the formulation of a scientific recipe for Cacio e pepe that capitalizes on our findings and highlights their applicative perspective.
@@
A potential future direction could be to better understand the starch-dependent morphology of the cheese clumps. As observed in Figs. 1 and 2, increasing the starch concentration causes the cheese-rich phase to abruptly switch from a large, system-spanning clump (the “Mozzarella Phase”) to many smaller aggregates. This phenomenon could be explained using a model that explicitly accounts for at least three distinct species. The simplest description would account for proteins in two states A and B with stronger and weaker interaction propensity, respectively, and an effective solvent encompassing water and all remaining molecules. Then one would need to account for transitions between the two states of the proteins, controlled by temperature and starch amount. In the spirit of Ref. [48], the mozzarella-like clump could represent a phase where A proteins are dominant, while small aggregates could arise whenever B proteins drive phase separation. From a microscopic point of view, A proteins could correspond to denatured whey proteins, while B proteins could encompass folded whey proteins and unfolded ones sequestered by the starch. One could even include casein micelles explicitly as a fourth component, instead of incorporating them in the effective solvent. An even more detailed model could consider protein aggregates of different sizes explicitly. One could then couple phase separation and gelation, describing the switch to a Mozzarella Phase as the onset of a gel phase [49, 50, 51]. Such generalizations of our model could shed more light on the molecular features of this system, eventually obtaining novel insights not only into perfecting the recipe but also in the broad field of food science. However, the main issue with these multi-component approaches is that they require more experiments to estimate the dependencies of all interaction and size parameters on temperature. For example, it would be necessary to quantify the fraction of denatured whey protein for each temperature in each phase, which is a hard task.
@@
Other interesting future directions could include a more in-depth analysis of how starch affects effective parameters and influences the viscosity of Cacio e pepe sauce, as well as the potential role of pepper grains, the other key ingredient of the recipe, as aggregation nuclei. We hope this paper has sparked the idea that a genuine passion for fine cuisine can be translated into insightful scientific investigations, refining complex preparations and making them more accessible with everyday kitchen tools.
Supplementary Material
The supplementary material contains: i) a detailed description of equipment, ingredients and experimental protocol employed in this study, along with a schematic representation and the actual picture of the experimental setup; ii) a discussion of the data analysis methodology, including three examples of segmented images used to quantify the degree of separation, along with a comparison of different possible measures for assessing the mixture’s phase behavior; iii) two analytical formulas representing the expressions of $\chi$ and $n$ of the model (see Eq. (2)).
@@ Acknowledgments
The authors thank Frank Jülicher, for supporting the initiative and stimulating discussions. We thank Tetsuya Spippayashi for enlightening clarifications on the historical origins of Cacio e pepe and Giuseppe Ricchitelli for helping with the construction of the experimental apparatus. We further thank Martina Gaiba, Alessandro Gaiba, John D. Treado, Virginia Lepore, Eleonora Nanu, Julia Kirsch, Lara Koehler, Burak Budanur, Irina Pi-Jaumà, Elizabeth Brückner, M.J. Franco Oñate, Giorgio Nicoletti, and Marco Salvalaglio for their support and for eating up the sample leftovers. Finally, we thank Simone Frau for taking the photograph in Fig. 1(a).
@@
Supplementary Material for
Phase behavior of Cacio e Pepe sauce
S1 Equipment, experimental protocol and ingredients
@@
    Sauce preparation: the ingredients for a small batch of sauce ($\approx 110$g, with the exact weight depending on the specific formulation) are mixed in a clean mug to achieve the desired composition of cacio cheese, starch, and water. The mixture is homogenized using an immersion blender, then transferred into a saucepan and weighed. The saucepan is then placed into the wooden platform of the heating apparatus (Fig. S1(a)).
@@
    Temperature ramp: the sauce is gradually heated up while its actual temperature is continuously monitored using a digital cooking thermometer with a probe immersed in the sauce itself. During the heating process, the mixture is constantly stirred with a spoon to ensure uniform heating and to prevent the formation of large cheese aggregates on the saucepan walls.
@@
    Image acquisition: the sauce sample withdrawn from the saucepan is quickly transferred into a clean Petri dish. The mixture is evenly spread across the dish by gently shaking it, and then the dish is positioned on the transparent support (Fig. S1(b)). A photograph of the sample is captured using the cell phone. After imaging, the sauce sample is collected in a separate container for later consumption. This entire operation is optimized to be completed within approximately 20 seconds, ensuring that no significant precipitation of cheese clumps occurs during the sample’s transfer and photography.
@@
Ingredients
The two ingredients employed in this work are dry corn starch, which was purchased in a common supermarket, and the pecorino cheese. For the latter, we opted for a Pecorino Romano DOP which has not to be confused with a common romano cheese and can be recognized by the logo in Fig. S3. The acronym DOP, which stands for "Denominazione di Origine Protetta" (protected designation of origin), is a European Union designation which indicates products that have been produced, processed and developed in a specific geographical area, using the recognized know-how of local producers and ingredients from the region concerned. This fact guarantees the quality of the cheese and its abiding to strict rules of production, in particular a defined list of ingredients. For our work we selected a Pecorino Romano DOP from Sardinia, Italy. Indeed, even if the name of the cheese indicates Lazio as the origin of the product, nowadays most of its production is done in the island of Sardinia. The complete nutritional information of the cheese used in our work is reported in Table S1.
Figure S3: Official logo of Pecorino Romano DOP and original DOP products
Energy 1643 kJ / 396 kcal
Fat 32 g
of which saturated 20 g
Carbohydrates < 0.50 g
of which sugars < 0.50 g
Protein 27 g
Salt 3.5 g
Table S1: Nutritional declaration of the cacio cheese employed in this work.
@@
The data analysis involved segmenting aggregates in petri dish images and computing their size. Initially, the images were converted to greyscale and normalized to an intensity range of $[0,1]$. To enhance visibility, we adjusted exposure using the equalize_adapthist function from the exposure module in scikit-image, with a clip limit of $0.01$. Images were then smoothed with a Gaussian filter using a sigma of $1$ pixel. These parameters were determined based on iterative trials.
@@
Figure S4: Three examples of segmented images illustrating varying degrees of separation: (a), (b) low, (c), (d) medium, and (e), (f) high. Panels (a), (c), and (e) show cropped grayscale experimental images (not to scale), where darker regions represent aggregates. Corresponding segmented aggregates are displayed in panels (b), (d), and (f), obtained using the histogram-based segmentation method described below.
@@
The segmented aggregates were then analyzed using the regionprops routine from scikit-image. In particular, we reported the mean aggregate size in the phase diagrams, calculated using the major axis length property from regionprops.
To compute the major axis length, which we use as a proxy for the aggregate size, one first calculates the covariance matrix $C_{ab}$ for the pixel coordinates $\{x_{i},y_{i}\}$ of each aggregate, using the following flat measure:
@@
with analogous formulas for $C_{xx}$ and $C_{yy}$. A covariance matrix carries the information of the aggregate area as $\sqrt{\det C}$. However, this area differs from the actual aggregate area, and so the matrix $C_{ab}$ was scaled such that $\det C=A^{2}$, where $A$ is the area of the segmented region. The largest eigenvalue of the rescaled $C_{ab}$ corresponds to the square of the aggregate’s size.
Figure S5: Visual guide to understand how an ellipse is used to measure an aggregate size. a) Segmented aggregate. b) Best ellipse overlaying the aggregate. c) Visual guide to identify the ellipse size via its major axis length.
In the main text, we used the major axis of the detected ellipses as a proxy for clump sizes. The average of this quantity over all clumps served as our order parameter, allowing us to assess the degree of separation in the mixture. For completeness, we compare phase diagrams constructed using different order parameters. In particular, in addition to the ellipses’ major axis (Fig. S6(a)), we also consider the area fraction (Fig. S6(b)) and a “rescaled size” (Fig. S6(c)). The area fraction is defined as the total area occupied by aggregates divided by the total sample area, properly restricted to the Petri dish region. The rescaled size, on the other hand, is obtained by multiplying the average major axis length of the ellipses by the area fraction. This measure was introduced to mitigate the impact of a single large clump dominating the statistics. Importantly, notice that all size measures lead to qualitatively similar phase diagrams.
@@
Figure S6: Quantification of the degree of separation of the data in Fig. 3 using different order parameters. In (a1)-(a3) the major axis of the ellipses averaged over the clumps is displayed, like in the main text. We compare it with the ellipse’s area (b1)-(b3) and the rescaled size, i.e. the product of area fraction times the average values of the ellipses’s major axes (c1)-(c3). For the different order parameters, we show the raw data (a1)-(c1), data after Gaussian filtering (a2)-(c2), and kernel regression smoothing (a3)-(c3).